# Betti numbers of the moduli space of Higgs bundles over a real curve

### v1

@@ Abstract.
We produce a formula for the $\mathbb{Z}_{2}$-Betti numbers of the moduli space $M_{r}^{d}$ of stable real Higgs bundles over a real projective curve, with coprime rank $r$ and degree $d$. Our approach relies on the motivic formula for the moduli space due to Mellit [M], Fedorov-Soibelman-Soibelman [FSS], and Schiffman [S], and the fact that the virtual $\mathbb{Z}_{2}$ Poincaré polynomial is a motivic measure over $\mathbb{R}$.
@@
Moduli spaces of Higgs bundles, symmetric products of a curve, spectral sequences, Betti numbers, anti-holomorphic involutions
@@ 1. Introduction
Let $\Sigma/k$ be a smooth, projective, geometrically connected curve of genus $g$ admitting a $k$-rational point, and denote by $M_{r}^{d}$ the moduli space of stable Higgs bundles of rank $r$ and degree $d$ over $\Sigma$. When $gcd(d,r)=1$, which we assume henceforth, $M_{r}^{d}$ is a smooth, quasi-projective variety.
@@
which is symmetric under permutations of the $\alpha_{i}$ and under $\alpha_{i}\mapsto q\alpha_{i}^{-1}$. If $k=\mathbb{F}_{q}$ and we set $\{\alpha_{i},\alpha_{i}^{-1}q|i=1,..,g\}$ are set to the eigenvalues of the Frobenius action on $H^{1}_{et}(\Sigma,\mathbb{Q}_{l})$ for $(q,l)=1$, then $A_{g,r,d}$ counts indecomposable bundles over rank $r$ and degree $d$ over $\Sigma$, and $q^{1+(g-1)r^{2}}A_{g,r,d}$ counts points in $M_{r}^{d}(\mathbb{F}_{q})$. Using a version of the Weil conjectures, it follows that that when $k=\mathbb{C}$, the Poincaré polynomial for the compactly supported rational Betti numbers of $M_{r}^{d}(\mathbb{C})$ equals
@@
Suppose that $\Sigma$ is a smooth, projective, geometrically connected curve curve of genus $g$ defined over $k=\mathbb{R}$, and suppose that $\Sigma(\mathbb{R})$ is a union of $b+1$ circles, with $b\geq 0$. Then
@@
where $\operatorname{Log}$ denotes the plethystic logarithm. Mellit proved that for all $r\geq 1$, $H_{g,r}$ is a Laurent polynomial in $q$, $z$ and $\alpha_{1},\ldots,\alpha_{g}$, and for all $d$,
@@
Our proof of Theorem 1.1 can be outlined as follows. Fedorov-Soibelman-Soibelman [FSS] proved a motivic version of Schiffman’s formula which, when $k$ is a field of characteristic zero, identifies the class $M_{r}^{d}]$ in the dimensional completion of the Grothendieck ring of Artin stacks over $k$, $\overline{Mot}(k)$. Mellit ([M] §6), identifies their class $M_{r}^{d}]$ as the image of $q^{(g-1)r^{2}+1}A_{r,d}$ under a homomorphism
@@
which sends the class $[X]$ of smooth, projective variety $X$ to the Poincaré polynomial $P_{t}^{vir}([X])=P_{t}(X(\mathbb{R}))$.
@@ 4. On purity of real semi-projective varieties
A semi-projective variety $X$ over $\mathbb{R}$ is a real variety equipped with an action of $\mathbb{R}^{\times}$ such that the base change $X\times_{R}\mathbb{C}$ is semi-projective over $\mathbb{C}$.
@@
Under these conditions $X^{G_{m}}=\coprod_{i\in I}F_{i}$ where the $F_{i}$ are non-singular, projective over $\mathbb{R}$, and we have a Bialycki-Birula stratification $X=\bigcup_{i\in I}U_{i}$ where
@@ 5. Average Betti numbers
This argument fails when $r$ is even. In this case $\Gamma_{r}$ intersects every connected component of $Pic^{0}(\mathbb{R})$, so (5.2) maps to only one of the $2^{b}$ components of $M_{r}^{d}(\mathbb{R})$. Indeed, the connected components of $M_{r}^{d}(\mathbb{R})$ have different Poincaré polynomials in general. For example (see [B] §5.2), when $g=b=2$, $M_{2}^{1}(\mathbb{R})$ has four components, three of which have Poincaré polynomial $t^{5}(1-t)^{2}(5t^{3}-7t^{2}+3t-1)$ and one of which has Poincaré polynomial $t^{5}(1-t)^{2}(t^{3}-3t^{2}+3t-1)$. Adding these together gives $4t^{5}(1-t)^{2}(4t^{3}-6t^{2}+3t-1)$. A priori, it is not at all obvious that the average value of the Poincaré polynomials of the connected components should have integer coefficients.

### v2

@@ Abstract.
We produce a formula for the $\mathbb{Z}_{2}$-Betti numbers of the moduli space $M_{r}^{d}$ of stable real Higgs bundles over a real projective curve, with coprime rank $r$ and degree $d$. Our approach relies on the motivic formula for the moduli space due to Mellit [M], Fedorov-Soibelman-Soibelman [FSS], and Schiffmann [S], and the fact that the virtual $\mathbb{Z}_{2}$ Poincaré polynomial is a motivic measure over $\mathbb{R}$.
@@
Moduli spaces of Higgs bundles, symmetric products of a curve, real algebraic geometry, motivic ring, Poincaré polynomials, Plethystic exponential
@@ 1. Introduction
Let $\Sigma/k$ be a smooth, projective, geometrically connected curve of genus $g$ over a field $k$ admitting a $k$-rational point, and denote by $M_{r}^{d}$ the moduli space of stable Higgs bundles of rank $r$ and degree $d$ over $\Sigma$. When $gcd(d,r)=1$, which we assume henceforth, $M_{r}^{d}$ is a smooth, quasi-projective variety.
@@
which is symmetric under permutations of the $\alpha_{i}$ and under $\alpha_{i}\mapsto q\alpha_{i}^{-1}$. If $k=\mathbb{F}_{q}$ and $\{\alpha_{i},\alpha_{i}^{-1}q|i=1,..,g\}$ are set to the eigenvalues of the Frobenius action on $H^{1}_{et}(\Sigma,\mathbb{Q}_{l})$ for $(q,l)=1$, then $A_{g,r,d}$ counts indecomposable bundles of rank $r$ and degree $d$ over $\Sigma$, and $q^{1+(g-1)r^{2}}A_{g,r,d}$ counts points in $M_{r}^{d}(\mathbb{F}_{q})$. Using the Artin-Grothendieck Comparision Theorem, it follows that that when $k=\mathbb{C}$, the Poincaré polynomial for the compactly supported rational Betti numbers of $M_{r}^{d}(\mathbb{C})$ equals
@@
Suppose that $\Sigma$ is a smooth, projective, geometrically connected curve of genus $g$ defined over $k=\mathbb{R}$, and suppose that $\Sigma(\mathbb{R})$ is a union of $b+1$ circles, with $b\geq 0$. Then
@@
where $\operatorname{Log}$ denotes the plethystic logarithm. Mellit proved that for all $r\geq 1$, $H_{g,r}$ is a Laurent polynomial in $q$, $z$, and $\alpha_{1},\ldots,\alpha_{g}$, and for all $d$,
@@
Our proof of Theorem 1.1 can be outlined as follows. Fedorov-Soibelman-Soibelman [FSS] proved a motivic version of Schiffmann’s formula which, when $k$ is a field of characteristic zero, identifies the class $M_{r}^{d}]$ in $\overline{Mot}(k)$, the dimensional completion of the Grothendieck ring of Artin stacks over $k$. Mellit ([M] §6), identifies their class $M_{r}^{d}]$ as the image of $q^{(g-1)r^{2}+1}A_{r,d}$ under a homomorphism
@@
which sends the class $[X]$ of a smooth, projective variety $X$ to the Poincaré polynomial $P_{t}^{vir}([X])=P_{t}(X(\mathbb{R}))$.
@@ 4. On purity of real semi-projective varieties
A semi-projective variety $X$ over $\mathbb{R}$ is a real variety equipped with an action of $\mathbb{R}^{\times}$ such that the base change $X\times_{R}\mathbb{C}$ is semi-projective over $\mathbb{C}$. When $\Sigma$ is defined over $\mathbb{R}$, $M_{r}^{d}$ is a semi-projective variety, with respect the $\mathbb{C}^{\times}$ action of scaling the Higgs field.
@@
Under these conditions $X^{\mathbb{C}^{\times}}=\coprod_{i\in I}F_{i}$ where the $F_{i}$ are non-singular, projective over $\mathbb{R}$, and we have a Bialycki-Birula stratification $X=\bigcup_{i\in I}U_{i}$ where
@@ 5. Average Betti numbers
This argument fails when $r$ is even. In this case $\Gamma_{r}(\mathbb{R})$ intersects every connected component of $Pic^{0}(\mathbb{R})$, so (5.2) maps to only one of the $2^{b}$ components of $M_{r}^{d}(\mathbb{R})$. Indeed, the connected components of $M_{r}^{d}(\mathbb{R})$ have different Poincaré polynomials in general. For example (see [B] §5.2), when $g=b=2$, $M_{2}^{1}(\mathbb{R})$ has four components, three of which have Poincaré polynomial $t^{5}(1-t)^{2}(5t^{3}-7t^{2}+3t-1)$ and one of which has Poincaré polynomial $t^{5}(1-t)^{2}(t^{3}-3t^{2}+3t-1)$. Adding these together gives $4t^{5}(1-t)^{2}(4t^{3}-6t^{2}+3t-1)$. A priori, it is not at all obvious that the average value of the Poincaré polynomials of the connected components should have integer coefficients.